OpenStudy (anonymous):

find midpoints of ab bc and ac graph and label points x y and z a(00) b(12,6) c(18,0)

OpenStudy (accessdenied):

Midpoints are calculated by averaging the x-values and y-values of the points: \( \displaystyle \text{Midpoint of } (x_1, y_1) \text{ and } (x_2, y_2) = \left( \frac{x_1 + x_2}{2} , \; \frac{y_1 + y_2}{2} \right) \)